The government has policy to make a highway as soon as possible. So, the constructor used the heavy machine instead of using man labour. A machine operator dug out a field in the shape of circular having radius 3 m and some depth 10 m. answer the following question.
A. When the machine operator dug out the field in the given dimension , the three dimensional shape of the digging area is ?
B. Find the volume of the soil .
C. How much cover area of the field dug out ?
D. Find the total inner surface area of the digging field.
E. If the cost of digging out of 1 cubic metre field is Rs 50 . find the total cost of complete digging out field.
1
Expert's answer
2020-12-16T20:03:44-0500

A. Cylinder


B. Volume of a cylinder ="\pi r^2h\\"

"=\frac{22}{7}\times 3^2 \times 10\\
=282.8571419m^3\\
\approx 282.857m^3"

C. The cover area of the field dug out is "\pi r^2". Since the cover is circular.

"=\frac{22}{7} \times 3^2\\
=28.2857m^2\\
\approx 28.29m^2"

D. Total inner surface area "=2\pi r(r+h)\\
=2.\frac{22}{7}.3(3+10)\\
=245.14286m^2\\
\approx 245.14m^2"

E. "1m^3=Rs50\\
282.857m^3=282.857\times Rs50\\
=Rs14142. 85"